given the linear equation y = 1/2x - 6 , find the y coordinates of the points
(-6,), (-2,) , and (4,), please show work. Plot those points and graph the linear equation.

Answer by ewatrrr(24785)


Hi
CHECKING our Answer***
y = 1/2x - 6
-9 = 1/2*-6 - 6 |Ordered pair Pt(-6,-9)
-7 = 1/2*-2 - 6 |Ordered pair Pt(-2,-7)
-4 = 1/2*4 - 6 |Ordered pair Pt(4,-4)